If (-5,3) is the endpoint of a line segment, and (-7,7) is its midpoint, find the other endpoint.
Answer by jim_thompson5910(35256)

(-5+x)/2 = -7

-5+x = -14

x = -9

------------------------

(3+y)/2 = 7

3+y = 14

y = 11


So the other endpoint is (-9, 11)